### v1

Abstract
With the advent of generative Artificial Intelligence (genAI), the notion of an agent has seen a resurgence in popularity. This has also led to speculation about the extent to which business process management, as a discipline and research field, may impact and be impacted by the deployment of genAI-based agents. To better ground such speculations into the state-of-the-art, we draw from the past 30 years of research on agents and business process management to establish the concept of Agentic Business Process Management (agentic BPM) that is only loosely coupled to the genAI hype. We conduct a series of interviews with BPM practitioners to explore their understanding, expectations, and concerns related to agent autonomy, adaptability, human collaboration, and governance in processes. The findings reflect both challenges with respect to data inconsistencies, manual interventions, identification of process bottlenecks, actionability of process improvements, as well as the opportunities of enhanced efficiency, predictive process insights and proactive decision-making support. While the technology offers potential benefits, practitioners also anticipate risks such as biases, over-reliance, lack of transparency, and job displacement within organizations. These concerns underscore the need for a robust methodological framework for managing agents in organizations.
Keywords:
autonomous agents generative AI business process management
1 Introduction
For over three decades, agents have periodically surged in Business Process Management (BPM), only to fade again. The 90s saw early excitement around goal-oriented software agents [20, 21], followed by the rise of Robotic Process Automation (RPA) in the late 2010s [12, 11, 17], promising efficiency gains for knowledge work. However, inflated expectations and high maintenance costs led to failures and disappointments in RPA adoption [19]. Now, Large Language Model (LLM)-based agents fuel another wave of optimism [9, 4], raising the question: what will remain once this hype subsides?
To foster lasting innovation, we propose a foundational perspective on agentic BPM. At its core, agentic BPM envisions autonomous software entities adapting to uncertainty rather than rigidly following predefined workflows. However, a key challenge remains: aligning individual agents (micro-level) with structured business processes (macro-level). While BPM seeks to orchestrate work efficiently, real-world processes often involve socio-technical nuances that rigid frameworks may overlook.
To explore these challenges, we conducted interviews with BPM practitioners, examining their views on agent autonomy, adaptability, human collaboration, and governance. Our findings highlight both opportunities, such as efficiency gains, predictive insights, and proactive decision making, and risks, including biases, over-reliance on automation, lack of transparency, and job displacement. These concerns underscore the need for a robust methodological framework for managing autonomous agents in organizations.
By reflecting on BPM’s evolving relationship with software agents and incorporating real-world perspectives, this paper aims to define agentic BPM as a sustainable, technology-agnostic approach to managing autonomous processes in organizations.
The rest of the paper is organized as follows. Section 2 reviews the history and evolution of agents in BPM. Section 3 defines agentic BPM and summarizes prior research on the interplay between agents for BPM and BPM for agents. Section 4 presents findings from a qualitative study with BPM practitioners on agentic AI. Sections 5 and  6 discusses the core pillars of agentic BPM and concludes its potential value for researchers and organizations.
2 Background
In this section, we provide a brief overview of the history and evolution of agents in BPM. We highlight both the continuous research progress in areas such as process mining and multi-agent reinforcement learning (Sect. 2.1), and hypes, most notably the rise and fall of RPA (Sect. 2.2). Finally, we discuss the current AI agent hype (in particular developments in conversational applications), relating to high-profile developments in industry (Sect. 2.3).
2.1 Beyond the Hypes: Agents in BPM
The integration of agent technologies for business process execution and their coordination began in the early 1990s. Fundamentally, an agent is a computer program that “operates autonomously, perceives the environment, persists over a prolonged time period, adapts to change, and creates and pursues goals” [34, pp.21–22] [^1] Definition of agents vary; discussing these definitions is out of the scope of this paper.. In the context of agents and Multi-Agent Systems (MAS) [^2] We may use the terms agents (plural) and MAS interchangeably, i.e., we consider a collection an agent necessarily an MAS., a business process can be seen as a collection of such agents that interact, given some orchestrated interdependencies, e.g., in a joint organization that imposes common goals [22, 20, 21]. Over the past decades, researchers have attempted to apply agents and MAS to business process management, with the objective to facilitate more dynamic execution behavior—often across organizations—while still maintaining high-level orchestration control [2]. However, initial approaches to agent-based business process management systems suffered from several disadvantages like lack of overall system control, as well as trust and delegation challenges [16].
On the process modeling side, related ideas have found their ways into the notion of process choreographies, representing the sequence and conditions under which multiple cooperating independent agents exchange messages in order to perform a task to achieve a goal state [8]. Indeed, choreographies, although arguably less known to the process modeling mainstream, are covered by the Business Process Model and Notation (BPMN) standard [32]. With the rise of blockchain technologies, the application of agents and business processes gained additional traction, at least in terms of potentially applied research: a well-functioning blockchain necessarily represents a network of autonomous agents, and blockchain technologies were, in the late 2010s, considered a crucial facilitator of future generation business process execution technologies [31]. For instance, choreography models were proposed as enablers of the model-driven implementation of data sharing and decision logic via smart contracts [28]. While blockchain and BPM can, clearly, be considered a (faded) hype, it did not have the notion of an autonomous agent at the center and neither made it into the BPM practice mainstream.
On the process analysis side, several agent-based approaches process mining to process simulation have been introduced [38, 37], most recently also to data-driven simulation that utilizes event logs [25]. Agent System Mining (ASM) combines process mining and ABM (Agent-based Modeling) to infer MAS models of operational business processes from real-world event data [40, 39, 36].
Fusing process analysis with execution, Reinforcement Learning (RL)-based approaches have been proposed as facilitators of adaptive decision-making through a trial-and-error approach [26, 35]: specifically, agents are employed to evaluate different process variants through RL-enhanced AB testing [35], refining their routing behavior based on performance feedback.
Finally, it is worth noting that a substantial body of work on agents exist that, while not explicitly connected to BPM, has substantial conceptual overlaps. Perhaps most notably, a long-running line of research on normative MAS [7, 14] studies how agents can and should act, in and across organizations, in accordance with norms and policies. These specify, somewhat analogous to process models or parts thereof, desired, expected, or enforced behaviors.
2.2 The first Hype: Robotic Process Automation
While agents and MAS address complex decision-making and optimization, RPA serves a different purpose. It is designed to efficiently automate simple, rule-based tasks on a large scale, which represent the majority of tasks in business processes [12].
Since 2015, research on RPA has highlighted numerous examples of how business process automation can significantly enhance performance [17]. When combined with BPM, RPA offers several advantages, including scalability, improved process accuracy, greater transparency and traceability, cost savings, and increased job satisfaction [12, 11]. However, successfully integrating RPA into practice remains a challenge, with 30-50% of initial projects failing during implementation [19].
The primary causes of RPA failures result from technical limitations. These include scalability challenges, the absence of error recognition mechanisms, inflexibility in UI integration, and data security risks that endanger sensitive business processes, making automation more complex. Additionally, incorrect cost estimation, as well as issues related to maintenance, governance, and reliance on human expertise, contribute to project failures [6, 11, 27].
Furthermore, employees, customers, and third parties may resist automation. Automated decision-making must comply with regulations, as the absence of human oversight can lead to legal and ethical risks. Automation may also create a black box effect, making it difficult for employees to understand the logic behind automated decisions [17]. Therefore, it is crucial to consider the effects of human-automation interaction alongside the technical aspects of introducing new technology [42].
Lastly, RPA is not suitable for full-scale business process automation, as it focuses on automating repetitive tasks rather than orchestrating end-to-end processes [27]. Since workflow automation and Integration platform as a service (iPaaS) are better suited for human-in-the-loop and API integration scenarios respectively, the automation use cases for RPA are reduced [43].
These limitations, combined with the absence of dynamic decision-making capabilities that ensure flexibility and adaptability, emphasize the need for an innovative process automation approach involving human-like intelligence and genAI-based agents [45].
2.3 The second Hype: genAI in BPM
With the rise of LLMs, BPM researchers, just like many other research communities, started to explore how the capabilities these new tools provide can be applied and integrated into BPM. In [41, 3, 5] all stages of the BPM lifecycle have been analyzed to identify use cases where LLMs can provide additional value to BPM. However, most of these papers focus on LLMs from a conversational perspective, leveraging them for tasks such as process understanding, information retrieval, and decision support through conversational interfaces. Additionally, current research and prototype implementations tend to cover only specific phases of BPM (such as discovery, analysis, and monitoring[10], rather than supporting BPM as an entire methodology. While one may claim that these approaches consider the LLM as an agent in the broader sense, they primarily focus on the automation of very specific tasks, leaving final decision-making, as well as broader goal-based autonomy primarily on the side of human users.
These strong limitations on software system autonomy are less prevalent in the long-term visions that have been developed against the backdrop of the generative AI (as well as other AI-based methods). In [9, 24, 33], AI-based systems in the broader sense are considered not only as applications that automate tasks within business processes and BPM, but also as intelligent systems beyond these applications. The difference between these systems and the previously mentioned concepts is that they are not only conversationally actionable and explainable but also have the potential for adaptability, self-improvement, and autonomy (to a deliberately limited, framed extent). More specifically, AI-Augmented BPM Systems (ABPMSs) leverage several AI-based methods to enhance process execution, analysis, and optimization while still operating within structured workflows. In this context, AI acts as a supporting mechanism, assisting human decision-making, process mining, and automation rather than fully controlling workflows [9]. Large Process Models (LPMs), on the other hand, follow a neuro-symbolic approach integrating pre-trained LLMs, potentially fine-tuned given extensive process data (e.g., process logs, BPMN diagrams, best practices) with more traditional symbolic systems that enable context-aware process design and analysis [24]. However, what the proposals have in common is that they primarily address the question of how AI-based technologies, both within and beyond the hype around LLMs, can facilitate business process management and execution.
What remains open is the question of how and to what extent BPM can best support the management of increasingly autonomous software agents that are deployed in organizations—and, consequently, what change in perspective is needed to provide such support.
3 Agentic BPM
From the literature review can follow that throughout the development of BPM, one of the main goals was (and still is) to achieve truly autonomous, end-to-end BPM systems capable of reasoning, planning, and dynamically adjusting processes without being confined to predefined workflows. Despite significant technological advancements, the reality remains that many systems still require substantial human involvement and intervention.
With the backdrop of the history of agents in BPM research, we set out to define Agentic Business Process Management. We align the definition with a set of desiderata, all reflecting the overall objective to frame agentic BPM in a sustainable manner that transcends technology trends and can thus stand the test of time.
Rooted in established terminology within BPM and AI.
Instead of redefining existing concepts or coining entirely new ones, the definition should primarily draw from key abstractions at the center of BPM and AI—or more specifically—autonomous agents and multiagent systems—research.
Technology-agnostic.
Acknowledging that technology trends tend to be either short-lived or result in the corresponding technology moving to the background as a commodity, the definition should not depend on any specific technology or groups/categories thereof.
Focused on BPM as a discipline, not on BPM software.
Our objective is to contribute to the study and practice of business process management, and not (merely) to improvements to BPM software.
Given these desiderata, we conceptualize Agentic Business Process Management as follows.
Concept 1 (Agentic Business Process Management)
eee
Agentic Business Process Management (ABPM) describes i) the deployment and execution of autonomous software agents in order to achieve business process goals, as well as ii) the application of agent-based abstractions for the process-oriented design and analysis of autonomous software agents.
Concept 1 makes use of several key concepts, all of which are well-established in the BPM and AI literature:
  • An (autonomous) software agent is a computer program that “operates autonomously, perceives the environment, persists over a prolonged time period, adapts to change, and creates and pursues goals” [34, pp.21].
  • A process goal operationalizes the objective that an organization strives to achieve with the corresponding business process, which—in turn–can be defined as “a set of activities that are performed in coordination in an organizational and technical environment”, to “jointly realize a business goal” [44].
  • An agent-based abstraction is a conceptual model explicitly featuring notions of (software or human) agents.
  • Process-oriented design and analysis describes the modeling business processes, as well as the drawing of inferences from data generated during process execution.
The clarifications above directly relate to well-known “text book” definitions of the key concepts of agents and business processes [^3] We omit common and nuanced discussions of what an agent is cf. [1, Chapter 1].. We do not assume agents are rational in the sense that they “act so as to achieve the best [expected] outcome” [34, pp.22]; with this, we acknowledge that agents may not always have explicit notions of goals, or may not work towards achieving individual or organization-level goals at all times. In contrast, we do emphasize that agents are software-based, i.e., we assume a context were self-directed work is not only executed by humans alone; this is the only technology commitment that our definition (deliberately) makes.
Finally, let us highlight that our definition attempts to cover key phases of the BPM lifecycle, by relating to the dichotomy of deployment and execution on the one hand, and design and analysis on the other (not distinguishing between discovery and analysis, steps that arguably tend to overlap in practice).
3.1 Agents for BPM versus BPM for Agents
Our proposed definition of ABPM is conceptual and not restricted to a particular technology in that it focuses on the application of BPM to MAS. Below, we further elaborate on the dichotomy of agents for BPM and BPM for agents, relating to several specific research lines and their key concepts in both the MAS and the BPM community. Figure 1 provides an overview of these concepts and their (approximate) emergence over time, separated by the dichotomy. Our ABPM concept concludes the BPM for Agents-line, thus establishing the proposal of a common denominator for earlier efforts.
Figure 1: Emergence of Abstractions and Technologies for Agentic Business Process Management over Time: Agents for BPM versus BPM for Agents.
Agents for BPM. Advancements in this category are technology-oriented and aim to automate or augment the execution of process management tasks. Depending on the nature of the technology, they can support BPM at different steps of the BPM life-cycle to different extents. Agent-Oriented Programming (AOP) focuses, in its application proposals to BPM, primarily on business process execution [20]). The same applies to RPA; in contrast, ABPMS are envisioned to support BPM holistically [9], and nascent research explores the application of LLM agents to process mining [4].
Agents for BPM can impact every phase of the BPM lifecycle, enabling better process control and process automation, while minimizing required user supervision (i.e., the extent of human involvement). For example, traditional agents and RPA can automate repetitive tasks with minimal human intervention, while genAI agents can offer real-time recommendations for optimizing processes. With LPMs, processes cannot only be created but also analyzed and predicted, and ABPMS can continuously evaluate process performance to drive process optimization and improve decision-making.
When these technologies work together, along with a clear outline of the final goals, they enhance process control, accelerate decision-making, and reduce the need for constant user supervision, ultimately driving greater automation and flexibility in BPM.
However, as we do not assume these technologies always operate with complete rationality (i.e., the concept of automation refers only to technologies being capable of executing and adapting business processes independently, based on their built-in capabilities), a clear framework is required to achieve meaningful BPM automation. This framework has to allow pragmatic and thoughtful utilization of appropriate technologies for specific tasks, alongside the strategic diversification of available resources.
BPM for Agents. It is crucial not only to integrate new agent-based technologies into BPM but also to adjust and extend BPM frameworks to support and handle the complexity of processes in organizations that deploy agent-based technologies. Since processes are often decentralized, dynamic, complex, and knowledge-intense, the organizations executing them can be viewed as MAS where humans—and increasingly also software systems—act with substantial degrees of autonomy.
Nonetheless, such processes, as well as the MAS executing them, need to be managed to align with organizational and broader societal requirements. From an MAS perspective, a long-running line of research on normative MAS studies this problem [7, 14]. It has some overlaps to BPM, e.g., in works that propose the application of deontic logic (formalizing notions such as obligation, prohibition and permission) to business process compliance [15]. Also, aspects such as the management and governance of agent autonomy is a problem that is of broad importance to the agent community, e.g., in the context of Web-scale systems [23], but has also been studied more specifically for business processes [29].
On the BPM side, the concept of process choreographies emerged in process modeling, outlining the sequence and conditions under which independent agents, working together, exchange messages to complete a task and reach a desired outcome [8]. Similarly, subject-oriented BPM shifts the focus from traditional process flows to subjects and their interactions, where subjects (e.g., human, software, or an agent) autonomously manage their tasks and decisions [13].
More recently, the need for Agent-Based Modeling and Simulation (ABMS) in BPM has been explored with increased attention [18]. Recognizing that processes often emerge from the interactions of autonomous agents, researchers have established the concept of Agent System Mining (ASM), i.e., the inference of agent-based models from event data [39].
Our proposed concept of ABPM encompasses these nascent attempts of viewing business processes as abstractions of somewhat agent-oriented socio-technical systems. Thus, our proposal advocates for the development of a systematic understanding of real-world MAS from a BPM perspective.
4 Practitioners’ Perspectives
In order to ground the concept of ABPM—as well as the “agents for BPM versus BPM for agents”-dichotomy—in real-world perspectives, we conducted a series of interviews with relevant BPM practitioners. Several process management software vendors, including UiPath, Salesforce, IBM, and Workday, use the term agentic AI to describe the interplay between genAI and agents. Thus, we leveraged the term during the interviews as a starting point to discuss the impact of agents on process management, expecting it to resonate effectively with the interviewees. Our study leverages a qualitative research design to examine the perspectives of professionals in business process management and automation, focusing on the integration and impact of agentic AI. It investigates the participants’ understanding, expectations, and concerns related to agentic AI’s autonomy, adaptability, human collaboration, and governance. Alongside discussing the characteristics of agentic AI, the study places greater emphasis on its impact on practitioners’ organizations. Therefore, the study also explores to what extent Agentic Business Process Management as a methodological framework is needed to address management needs that arise from greater software autonomy.
4.1 Interview Methodology
To systematically examine the perceptions and attitudes towards agentic AI, a qualitative content analysis approach [30] was followed. Semi-structured interviews were conducted with professionals across various industries, focusing on their understanding, expectations, and concerns regarding autonomy, adaptability, human collaboration, and governance of agentic AI. The process began with transcription, familiarization with the interview data and analysis using a deductive-inductive coding approach. Initial deductive categories were derived from the research questions, while additional inductive subcategories emerged from the data, allowing themes to derive directly from the participants’ responses. A total of 22 participants from various industries and roles were interviewed (see Fig. 2). Sessions lasted 60 to 90 minutes, with some conducted in pairs or groups due to accessibility constraints.
Figure 2: Demographic Distribution across Participants
4.2 Results
The interview results highlight various themes mentioned by the interviewees, which are discussed below.
Understanding. Among the 22 participants, 10 were familiar with the term agentic AI, describing it as a self-learning technology that operates autonomously and adapts to its environment. Some viewed it as an evolution of RPA, overcoming technical limitations with AI. Others associated it with a digital assistant or an orchestration layer that coordinates tasks across specialized agents. In this context, a participant highlights:
”[Agentic AI is an agent] that can do significantly more than dumb botting, where I always have to tell it everything exactly and have precisely structured it, and it really only helps where I have very clearly structured, stable processes.”
Importantly, none of the participants reported to have actual practical experience with agentic AI, i.e., the assessments participants provide is based on expectations, often against the backdrop of existing agent-like automation technologies such as RPA.
Benefits. Participants highlighted several potential benefits of integrating agentic AI into organizational processes (see Figure 3 (a)). The technology is expected to enhance efficiency by automating routine tasks, streamlining operations, and reducing errors, allowing employees to focus on higher-value work and achieving time and cost savings. It is also anticipated to improve data quality, ensuring consistent and accurate handling of information. Better compliance was another potential benefit, as agentic AI could monitor regulations and enforce standards automatically. Scalability was frequently noted, enabling businesses to handle larger workloads without proportional increases in staffing. Additionally, it is believed to democratize process data, making it more accessible and usable. In this context, a participant highlights:
”More employees could initiate changes or optimizations if the software supports them directly, reducing dependency on specialized roles.”
Risks. Despite its potential, implementing agentic AI carries risks (see Figure 3 (b)), such as bias from flawed training data, over-reliance leading to diminished human judgment, and lack of transparency in AI-driven decisions. Participants also raised concerns about cybersecurity threats, job displacement, and unauthorized decision-making, stressing the need for adequate human oversight to prevent unintended consequences. In this context, a participant highlights the cultural aspect:
”It’s a cultural thing to be able to accept autonomy and decision making being taken away from a human.”
Figure 3: Agentic AI among Participants: Expected Benefits and Risks.
Use cases. Participants identified several potential use cases for agentic AI (see Fig. 4 (a)). Key applications include process monitoring to detect inefficiencies and suggest improvements, and predictive analytics to forecast trends and provide actionable insights. Task automation was another major focus, with agentic AI handling routine activities like data entry and document processing. Specific tasks include master data maintenance, user administration, root cause analysis, and decision support through dashboards. A participant highlights:
”Processes often get stuck due to errors in master data, such as mismatched product codes or pricing issues. [Agentic] AI could analyze and fix these autonomously.”
Agentic AI could enhance customer service, while supply chain optimization benefits from agentic AI managing inventory and predicting demand. In finance, agentic AI could aid in fraud detection and transaction monitoring. Additionally, agentic AI could structure unstructured datasets, improving accessibility and analysis.
Requirements. Participants stressed the need for clear rules and guidelines to ensure agentic AI operates ethically and transparently. They highlighted the importance of audit logs, data retention, transparency, robust security, and adherence to corporate policies and regulations. In this context, a participant highlights:
”[The agent] would basically replace an FTE [^4] Full-Time Equivalent, i.e., a common unit to measure work in terms of person efforts., let’s just put it that way; you also have to provide it with the same framework that the employee would be confronted with because what would the employee do if they encounter difficulties?”
Defining roles, responsibilities, and limitations for agentic AI is crucial, along with compliance with data protection laws and risk management frameworks. Seamless integration with existing processes prevents disruptions, while comprehensive employee training ensures effective use of agentic AI. Managing costs, including setup, maintenance, and upgrades, was also highlighted as essential. Preconfigured use cases and applications were recommended to build trust by demonstrating agentic AI’s capabilities and delivering clear benefits. Figure 4 (b) gives an overview of requirements that practitioners highlighted.
Figure 4: Agentic AI among Participants: Use Cases and Roll-out Requirements.
Autonomy and adaptability. Participants stressed the importance of managing agentic AI’s autonomy to maintain accountability, reliability, and safety. While agentic AI can independently handle low-risk tasks like data preparation, human oversight is essential for critical decisions or significant changes. Configurable autonomy was recommended, granting agentic AI flexibility in low-risk areas while restricting it in high-risk scenarios, such as system changes or external communications. A gradual approach was also suggested, starting with simple tasks and increasing autonomy as trust in the system grows. In this regard, a participant highlights:
”It shouldn’t make changes in source systems or install new apps autonomously. That crosses the line because those areas are managed by different teams and require coordination.”
Adaptability is a key strength of agentic AI when operating within clear boundaries, but concerns about trust and consistency highlight the need for human oversight to monitor and guide its adaptations. Transparency was highlighted as essential, with participants calling for clear documentation of all AI decisions and actions. In this context, a participant mentions:
”If it’s routine, the decision itself should be documented. The more complex the task, the more I want to see how the process was developed.”
This would include high-level summaries for management and detailed logs for technical teams to ensure traceability and allow for changes to be reverted if necessary. Participants acknowledged agentic AI’s potential to enhance workflows and efficiency but emphasized the need to operate within predefined rules to prevent unintended outcomes. Agents should support, not replace, humans in tasks requiring complex judgment. Human validation and oversight remain crucial for critical scenarios.
Human involvement and responsibility. Participants agreed that although agentic AI can autonomously handle low-stakes decisions, it should serve as a decision-support tool for complex process scenarios, ensuring human involvement to maintain accuracy and accountability. Its role is to provide clear analyses, suggest actions, and outline impacts, enabling well-informed human decisions. Some proposed methods like briefings or notifications to engage humans in urgent situations, viewing agentic AI as a collaborative assistant offering insights and recommendations. In this regard, a participant underlines:
”For complex decisions, the [agentic] AI should provide full context; how it arrived at the decision, what data it used, and the potential consequences; just like how humans consult colleagues for advice.”
Responsibility for agentic AI failures was seen as a shared effort among organizations, developers, and business leaders. Organizations deploying agentic AI bear ultimate responsibility, particularly by providing oversight in critical tasks to validate agentic AI decisions. Developers play a key role in building reliable systems, while organizations establish robust oversight mechanisms. Key users and application owners were identified as primary contacts for resolving issues, with some suggesting dedicated teams to manage and ensure accountability for agentic AI. Figure 5 summarizes responsibilities of agentic AI failure that practitioners highlighted.
Figure 5: Agentic AI among Participants: Responsibility for Failures.
5 Discussion
The findings show that organizations need a clear management framework, which we define as Agentic Business Process Management, to successfully introduce agents into their business processes. This framework aims to help organizations adopt agents effectively while ensuring humans stay informed and involved. The interview results underscore the two main pillars of agentic BPM. First, it links business context, guardrails, and human-agent collaboration to the deployment and execution of agents, ensuring these technologies help achieve process goals.
Business context.
Organizations should define clear business goals for agents, specify the expected benefits, and estimate the costs associated with its introduction and maintenance.
Guardrails.
Organizations must ensure that agents operate within the boundaries of legal, ethical, and organizational rules, while considering the context and environment in which the agents function.
Human-agent collaboration.
Clear roles and responsibilities should be established to enable effective collaboration between humans and software agents, including identifying situations where human intervention is necessary in case of agent failure.
Second, it ties customization, risk management, and adoption, to the application of agent-based concepts for effectively designing and analyzing processes involving autonomous software agents.
Customization.
Organizations should tailor agents to meet their specific needs, determining factors such as the level of autonomy granted to agents, how human oversight and review are conducted, the extent of documentation required, and the desired level of reasoning and transparency in agent operations.
Risk management.
Organizations must implement safety measures to monitor performance and address potential risks, including preventing undesired adaptability or evolution of the agents.
Adoption.
Organizations must ensure the seamless and responsible integration of agents into existing business processes and systems, while implementing fallback mechanisms to enable a safe reversion if necessary.
6 Conclusion
The rise of generative AI has led researchers and software vendors to explore its integration with agents, creating the concept of agentic AI. However, the impact of this technology and its management in critical processes remains unclear. Therefore, this paper set out to explore how BPM for agents—a methodological concept we define as agentic BPM—can assist organizations in addressing these challenges. Interviews of BPM practitioners highlight that the emerging generation of software agents may help automate repetitive tasks and enable humans to focus on meaningful work but also underline risks like bias, over-reliance, job loss, and lack of transparency. This emphasizes the importance of developing a methodology that ensures ethical use of agents through strong governance and active human oversight. The initial qualitative study presented in this paper highlights the need for broader research. Future studies should include participants with varied backgrounds, monitor the actual adoption of new generations of software agents (such as LLM-based agents), explore early adopters’ use cases, and assess the value of a defined agentic BPM framework. It should also investigate the application of agents in specific industries and develop strategies for its responsible adoption, balancing expected performance benefits with requirements regarding accountability and long-term system maintainability.